Two-fifths of a school’s 640 students eat a school lunch each day. How many students eat school lunches each day?
Found 2 solutions by greenestamps, Alan3354:
Answer by greenestamps(13200)



"Two-fifths of a school’s 640 students..."

The "of" means multiply, so

ANSWER: (2/5)(640) = ...

(you can do the calculation)


Answer by Alan3354(69443)

Two-fifths of a school’s 640 students eat a school lunch each day. How many students eat school lunches each day?
---------------
(2/5)*640
= 2*640/5
= 1280/5
etc
=============
Or,
= 0.4*640